### v1

# Large Language Model-driven Meta-structure Discovery in Heterogeneous Information Network

Abstract.
Heterogeneous information networks (HIN) have gained increasing popularity in the recent decade for their ability to capture the complex relations between nodes of diverse types. Meta-structure was proposed to identify the important patterns of relations on HIN, which has been proven effective for extracting rich semantic information from HIN and facilitating graph neural networks to learn expressive representations. However, hand-crafted meta-structures pose challenges for scaling up to large and complex HINs, which draw wide research attention for developing automatic meta-structure search algorithms. Previous efforts predominantly concentrate on searching for meta-structures that have good empirical prediction performance, overlooking the explainability of discovered meta-structures. As a result, they often produce complex meta-structures that are prone to overfitting and incomprehensible to human researchers. To address this challenge, we draw inspiration from the emergent reasoning abilities of large language models (LLMs). In this paper, we propose a novel reasoning meta-structure search (ReStruct) framework that integrates LLM reasoning into the evolutionary procedure for meta-structure search. ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate the semantically feasible meta-structures. Besides, ReStruct also employs performance-oriented evolutionary operations. These two competing forces allow ReStruct to jointly optimize for semantic explainability and empirical performance of meta-structures. Finally, we also design a differential LLM explainer that can produce natural language explanations for the discovered meta-structures, and refine the explanation by reasoning through the search history. Experiments on five representative HIN datasets demonstrate that ReStruct achieve state-of-the-art performance in both node classification and link recommendation tasks. Additionally, a survey study involving $73$ graduate students shows that the meta-structures and natural language explanations generated by ReStruct are substantially more comprehensible.
Heterogeneous Information Networks, Large Language Models, Graph Neural Networks.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Heterogeneous information networks (HINs) have been proven effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016). As a result, HINs are increasingly adopted in a wide range of applications, including social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020). To fully exploit the rich semantic information encoded in HINs, researchers proposed to use hand-crafted templates of relation sequences to model the complex proximities on HINs (Sun et al., 2011), which is known as meta-path. It was later extended to meta-structure to capture more general interaction patterns than linear paths. The recent advancement of heterogeneous GNNs showed that meta-structures can be leveraged to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019), which are crucial for various downstream applications. However, the previous works on HINs modeling predominated leverage hand-crafted meta-structures, which heavily rely on the knowledge of domain experts. Therefore, they are difficult to scale up to larger and more complex HINs that are vital for real-world applications.
Motivated by the importance of domain adaptive meta-structure design, numerous recent research efforts are dedicated to designing algorithms for automatic meta-structure design. Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search model (Ding et al., 2021) to automatically identify meta-structures that could lead to superior performance in heterogeneous GNNs. However, these previous attempts solely focus on the prediction performance of meta-structures, which often lead to highly complex structures that are difficult to explain and susceptible to overfitting. Such “meta-structures” deviate from the original inspiration of meta-structures research that aims to extract semantically clear features from HINs (Huang et al., 2016).
The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to address the challenges of meta-structure discovery. The scaled-up versions of LLMs show emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a). For example, researchers find chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b). Such general-purpose reasoning capability holds huge potential to make sense of the rich semantic information in HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms. That is, instead of searching for best-performing meta-structures, the reasoning power of LLM could facilitate the joint optimization for semantic explainability by emulating human reasoners. Such reasoning process could effectively prevent the overfitting of identified meta-structures, and produce human understandable knowledge from given HINs.
In this paper, we propose a novel ReStruct (short for Reasoning meta-Structure search) framework that integrates LLM reasoning into an evolutionary procedure for meta-structure search. Specifically, ReStruct employs a grammar translator to encode meta-structures with nested clauses in natural language (see Figure 2). It ensures the semantic meaning of meta-structures can be readily comprehended by LLMs as natural language sentences. Besides, we define a set of basic operations that can explore the feasible adjacent meta-structures of a given meta-structure, allowing ReStruct to search the adjacent possibilities of promising candidates. Different from pure performance-oriented search, ReStruct jointly evaluates the semantic feasibility and empirical performance to determine promising candidates. On one hand, ReStruct employs a few-shot LLM predictor to evaluate the semantic similarities between given meta-structure candidates and the searched meta-structures in past experience, and use a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities. On the other hand, after the semantically feasible candidates are empirically evaluated with heterogeneous GNNs, ReStruct uses a evolutionary updater adopting the classic elimination-reproduction procedure to mimic natural selection processes that refine meta-structure candidates based on their performances. Finally, we design a differential LLM explainer that generates natural language explanations for the discovered meta-structure. It employs a chain-of-thought prompting technique to perform structural comprehension and performance attribution step-by-step. Such a reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent, not chosen ones.
We evaluate ReStruct on five representative HIN datasets. Experiments show that ReStruct achieve state-of-the-art performance on both node classification and link recommendation tasks. Besides, ReStruct can generate natural language explanations as it searches through the solution space. To effectively evaluate the explainability of discovered meta-structures, we perform a user survey on $73$ graduate students with domain knowledge of HIN research. Survey results show that the meta-structure discovered by our ReStruct method is chosen as the most comprehensible one by 46.6% participants against 3 strong baselines, outperforming the second best baseline by 61.8%. More importantly, the natural language explanations generated by our method are preferred by most human participants (77.6% on average) in a head-to-head comparison with baseline methods.
We summarize our main contributions below:
  • We propose a novel ReStruct framework that integrates LLM reasoning into an evolutionary meta-structure search procedure. ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, which coordinates the competing forces of empirical performance-oriented evolutionary updater and semantic similarity-oriented LLM selector. ReStruct represents a path-breaking development in meta-structure search algorithms, which could generate meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
  • We design a grammar translator to encode meta-structures as natural language sentences, which unleashes the reasoning power of LLMs to make sense of the rich semantic information on HINs. On top of this, we design a differential LLM explainer that can generate human comprehensible natural language explanations for discovered meta-structures.
  • Extensive experiments show that ReStruct achieves state-of-the-art performance on five representative datasets. More importantly, a user survey validates that ReStruct substantially outperforms baseline methods in terms of the explainability of discovered meta-structures.
2. Preliminaries
Here, we provide the definitions of heterogeneous information networks, meta-paths, and meta-structures as in the literature.
Definition 3.1. Heterogeneous Information Network (HIN) (Sun et al., 2011). An information network (IN) is mathematically a graph denoted as $G=\{V,E,T,R,\sigma,\phi\}$, with $V=\{v_{1},v_{2},...,v_{n}\}$ being the set of nodes, $E=\{e_{1},e_{2},...,e_{m}\}$ being the set of edges, $T=\{t_{1},t_{2},...,t_{k}\}$ being the set of node types, and $R=\{r_{1},r_{2},...,r_{j}\}$ being the set of edge types. $\sigma:V\rightarrow T$ is a function that maps each node to its associated type, and $\phi:E\rightarrow R$ is a function that maps each edge to its associated type. The network schema of $G$ is then denoted as $S=\{T,R\}$. If $|T|>1$ (multiple types of nodes) or $|R|>1$ (multiple types of edges), $G$ is a heterogeneous information network (HIN). Otherwise, it is a homogeneous information network.
Definition 3.2. Meta-path (Sun et al., 2011). Given an HIN $G$, a meta-path $P=t_{1}\xrightarrow{e_{1}}t_{2}...\xrightarrow{e_{p-1}}t_{p}$, is a sequence of node types and edge types defined on the network schema $S$, connecting a single source node type and a single target node type. One meta-path may correspond to many meta-path instances in $G$.
Definition 3.3. Meta-structure (Huang et al., 2016). Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures between the source node type and the target node type.
3. Methods
In this section, we introduce our proposed methods in detail. In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLM’s understanding of its semantic meanings. In Section 3.2, we design three basic operations for generating candidate meta-structures. In Section 3.3, we design two LLM modules to evaluate and select candidate meta-structures with semantic similarity orientation. In Section 3.4, we combine LLM-guided optimization with evolutionary processes to form an effective derivative-free optimization framework. Figure 1 shows the overview of our framework.
Figure 1. Overview of our proposed ReStruct framework.
3.1. Natural Language Encoding of Meta-Structures
To facilitate the LLM’s comprehension, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2. For any meta-structure, we first traverse the structure to find all possible simple paths from the source node to the target node. Each resulting path is equivalent to a meta-path connecting the source and the target, decomposed from the original meta-structure. For each path, we encode it into a natural language sentence using nested clauses signified by THAT, which is a commonly-used grammar in English and thus hypothesized to be well-comprehended by the LLM. After obtaining the natural language encodings of the decomposed meta-path, i.e., sub-logics, we further combine them using a conjunction word AND to indicate the logical summation effect.
Figure 2. Natural language encoding of meta-structures.
3.2. Basic Operations for Candidate Meta-Structure Generation
To generate comprehensive candidates for LLM selection and meanwhile ensure validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing Lego. Examples are shown in Figure 3.
  • INSERTION. This operation replaces one edge of the original meta-structure with a component.
  • GRAFTING. This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively.
  • DELETION. This operation deletes a certain amount of nodes from the original meta-structure, and tries connecting all neighbors back.
In our experiments, we take all meta-paths with no more than $2$ nodes as components for INSERTION, and all meta-paths with no more than $3$ nodes as components for GRAFTING. DELETION does not require components as input, as it regards all existing nodes on the original meta-structure as operation candidates. Note that using components with more nodes can bring larger exploration space at each step, but may also result in chaos and confusion for the LLM, and thus we leave it as future work to investigate the optimal component settings.
Figure 3. Basic operations of exploring adjacent meta-structures.
3.3. Semantic Similarity-oriented LLM modules for Candidate Selection
To regulate the explainability of discovered meta-structures, we design two LLM modules that evaluate and select candidate meta-structures in a semantic similarity-oriented manner. We illustrate the interaction processes between the main program and the LLM agents in Figure 4.
3.3.1. Few-shot LLM Predictor
Following the execution of the aforementioned basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates. This collection comprises meta-structures that have been previously encountered and evaluated in earlier generations, as well as entirely novel ones. To inform the decision-making process with insights gleaned from previous evaluations, we design a few-shot LLM predictor to predict the performance $\hat{p}$ of each candidate based on a performance pool that stores the structures and corresponding performances of seen meta-structures. Additionally, the few-shot LLM predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate. Intuitively, if the few-shot LLM predictor deems a candidate highly similar to a counterpart in the performance pool, it is inclined to assign higher confidence to this prediction. This is grounded in the notion that structural similarity is likely indicative of functional similarity. An illustrative prompt-response round is exemplified in Step 1 of Figure 4.
3.3.2. Similarity-oriented LLM Selector
Receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design an Similarity-oriented LLM selector to make the final decision of picking one single candidate to enter the next generation. During this process, the Similarity-oriented LLM selector is expected to jointly consider, and perhaps trade-off between, a meta-structure’s semantic meanings, the expected outcomes of choosing it, as well as the credibility of this outcome expectation. An illustrative prompt-response round is exemplified in Step 2 of Figure 4.
3.4. Performance-oriented Evolutionary Updater
With closed-source LLM modules in the loop, it is not feasible for us to obtain the gradient for optimizing meta-structure search. Therefore, we operationalize a derivative-free optimization framework with inspirations from the genetic algorithm. Specifically, we maintain a population of $N$ individuals, each representing a distinct meta-structure. In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task such as recommendation and node classification. After evaluation, the underperforming meta-structures are eliminated from the population. The surviving meta-structures undergo a reproduction phase, where duplication occurs with probabilities proportional to their performances. In essence, this phase uses promising meta-structures to replenish the population to its original size. Both elimination and reproduction processes mirror natural selection mechanisms that enable species evolution in the wild. After getting the modified population, we feed it into the LLM to improve each individual meta-structure. Under the evolutionary framework, this can be regarded as a way of targeted mutation, as in this process new nodes and/or edges can be generated and some of the existing nodes and/or edges will be removed. The modified population will be further evaluated at the onset of the next generation, forming a loop of derivative-free optimization.
3.5. Differential LLM Explainer
One prominent advantage of the LLM lies in its unparalleled ability for natural language generation. To harness this ability, we design a differential LLM explainer module to automatically generate human-comprehensible textual explanations elucidating the reasons behind the superior performance of discovered meta-structures. We design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) to guide the differential LLM explainer in discerning the critical structural properties that contribute to performance enhancement. Specifically, the prompting process for analyzing a given meta-structure $T$ unfolds in the following two steps:
Step 1: Structural Comprehension. We first sample a set of $n$ one-step neighbors for $T$, and translate each of them into a natural language sentence according to Method 3.1. Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both the original meta-structure and all its sampled neighbors. This involves breaking down each of them into meaningful sub-structures and identifying the sub-structure functions.
Step 2: Performance Attribution. We first perform a quick evaluation of $T$ and each of the sampled neighbors separately by training a GNN with one structure at a time for downstream tasks. Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in $T$. This involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
Combination of these two steps empowers the differential LLM explainer to unravel the intricate connections between structural properties and performance, providing a comprehensive understanding of the discovered meta-structures. The effectiveness of this module is further justified by a pilot study involving human evaluators, which is elaborated in Section 4.4.2.
Figure 4. Example of LLM prompts and feedbacks.
4. Experiments
4.1. Experimental Settings
4.1.1. Datasets
We evaluate our proposed framework ReStruct on two important tasks in heterogeneous information network learning: recommendation and node classification.
For the recommendation task, the goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., businesses or movies). Under this task, we conduct experiments on three widely used real-world datasets: Amazon, Yelp, and Douban  [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding. Each dataset includes user ratings of items, which are converted to 0-1 binary labels with a threshold of $2$ to indicate the presence or absence of preference. Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into train-validation-test sets with a ratio of 3:1:1. The other half is reserved for network construction so as to prevent label leakage. We take all user-item pairs with the label ’0’ as negative pairs, and also randomly split them into train-validation-test sets to pair each positive pair. If the total number of negative pairs is not enough, we randomly sample unconnected items until reaching the desired number. We adopt AUC (the area under the ROC curve) as the evaluation metric.
For the node classification task, the goal is to predict the correct labels of nodes of a given type, such as the genre of a movie. Under this task, we conduct experiments on two other widely adopted real-world datasets: ACM and IMDB. In these datasets, the classification targeted node type is papers (subjects) for ACM and movies (genres) for IMDB, respectively. We follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019). We adopt the Macro-F1 score as the evaluation metric.
Details of all five datasets can be found in Appendix A.
4.1.2. Baselines
We compare ReStruct with a set of state-of-the-art baselines. These baselines can be classified into three categories:
  • Hand-crafted meta-paths: HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism.
  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework.
  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which searches meta-structures for recommendation on HINs with a genetic algorithm; (2) DiffMG (Ding et al., 2021), which formalized the problem into a neural architecture search flavor and search for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes with multi-graph search.
4.1.3. Hyperparameter Settings
For our model, we run the algorithm for $30$ generations with a population size of $5$ and an elimination rate of $0.2$. When modifying each meta-structure, we randomly sample a set of $20$ candidates if there are too many of them from the one-step neighbors. When predicting meta-structure performances with the LLM predictor, we randomly sample a set of $30$ evaluation records from the performance pool to enable a few-shot learning paradigm through prompting with examples. For the LLM, we use the GPT-4 model by calling the OpenAI API. For all baseline models, we follow the implementation released by the authors. We fix the hidden dimensions to $64$ for all evaluated models, and tune other hyperparameters including learning rate, weight decay, and dropout by referring to the performances on the validation set. To reduce the noise brought by randomness during program execution, for each combination of (model, task, dataset), we run experiments with $10$ different random seeds and report the average performance with standard deviation.
4.2. Comparison on Recommendation
In Table 1, we report the experimental results of our proposed framework on the recommendation task compared to baselines. First, we observe that models using meta-paths (HAN, RMSHRec) exhibit substantially poorer performances than those using meta-structures (all the rest), confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning. Second, ReStruct consistently achieves the best performance across three datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs. Specifically, the performance gain over GEMS confirms that LLM-guided “targeted mutation” converges to better solutions that pure random mutation in a classic genetic framework.
HAN RMSHRec GEMS DiffMG PMMM ReStruct
Amazon 63.53$\pm$0.24 61.80$\pm$0.15 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
Yelp 74.54$\pm$0.09 59.39$\pm$0.81 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
Douban 85.12$\pm$0.12 79.72$\pm$0.32 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
Table 1. AUC (%) of recommendation on three datasets.
Figure 5. Discovered meta-structures on Yelp and the corresponding generated natural language explanations.
4.3. Comparison on Node Classification
In Table 2, we report the experimental results of our proposed framework on the node classification task compared to baselines. First, HAN and RMSHRec do not achieve desirable performances. Specifically, HAN is limited by its heavy reliance on hand-crafted meta-structures, while RMSHRec suffers from high instability presumably due to the limitation accompanied by the reinforcement learning framework. Second, DiffMG and PMMM demonstrate improved performances, showcasing the advantage brought by meta-structures over meta-paths, as well as the NAS searching framework. Finally, the meta-structure discovered by ReStruct achieves the best performance on both ACM and IMDB datasets.
HAN RMSHRec DiffMG PMMM ReStruct
ACM 91.20$\pm$0.25 90.32$\pm$0.74 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
IMDB 55.09$\pm$0.67 56.72$\pm$2.60 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
Table 2. Macro F1 scores (%) of node classification on two datasets.
4.4. Explainability Analysis of Discovered Meta-Structures
4.4.1. Visualization of Discovered Structures and LLM-generated Explanation
In Figure 5, we present $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for the recommendation task, along with a summary text explaining the structural attributes underlying their outstanding performances, which are essentially output from our differential LLM explainer further condensed for clarity. We highlight the LLM-identified good (relevant) sub-structures in red, while the bad (distracting) sub-structures in blue. This visualization facilitates a comprehensive understanding of each structure’s composition. For example, the meta-structure in Figure 5 (a) contains critical yet simple sub-structures describing the geographical, social, and business category contexts of user behavior. While more nodes and edges can create complex relationships, they are not always desirable for HIN learning, as showcased by the identified distracting sub-structures. These structures, though challenging to handcraft, carry semantically meaningful explanations that remain accessible with textual assistance.
4.4.2. Pilot Study of Human Evaluation
We conduct a pilot study with $73$ human participants, more specifically, graduate students with domain knowledge of HIN research, to evaluate the level of explainability provided by our framework compared to baselines. We structure the study around two sets of questions. The first set aims to compare the explainability of the best-performing meta-structure discovered by our model with those found by several baseline models, including GEMS, DiffMG, and PMMM. Participants are asked to select one meta-structure from a set of meta-structures, each being the best one found by a specific model, as documented in the original papers (Han et al., 2020; Ding et al., 2021; Li et al., 2023). The second set of questions investigates the comprehension gain brought by the textual explanation generated by our differential LLM explainer, as described in Method 3.5, coined the Differential Explanation. As a baseline, we include the textual explanation generated by directly prompting an LLM to explain the reasons behind a given meta-structure’s strong performance, without undergoing the two-step prompting process, coined the Non-Differential Explanation. For three meta-structures among the best ones discovered by our model on the Yelp HIN, participants are presented with both types of explanations. They are then asked to determine which explanation is more helpful in enhancing their (1) understanding of how this meta-structure is constructed, and (2) insights on how to design a better one. The complete questionnaire utilized in our study is available in Appendix B.
Figure 6 shows the result of our first question set. Among all four models, the meta-structure discovered by our model is regarded by most people ($46.6\%$) as the most informative or comprehensible one, outperforming the second baseline, DiffMG ($28.8\%$), by $61.8\%$. GEMS and PMMM follow with the same level of recognition, $12.3\%$. This outcome aligns with our initial objective of discovering meta-structures that not only excel in downstream tasks but are also accessible and easily understood by human users, presumably because our framework seeks to find meta-structures that are semantically meaningful while keeping as simple as possible.
Figure 7 shows the result of our second question set. When comparing two types of generated textual explanations, the Differential Explanation is consistently and significantly more preferred by human participants ($77.6\pm 2.8\%$) across various meta-structures. This outcome justifies the effective design of our differential LLM explainer that unleashes the LLM’s reasoning ability on the intricate connections between sub-structures, sub-functions, and how they interact to determine the ultimate performances.
Figure 6. Human evaluated explainability of the best meta-structures discovered by different models.
Figure 7. Explainability gain brought by LLM-generated differential meta-structure explanation compared to non-differential explanation.
4.5. Ablation Study
To assess the efficacy of the designed grammar translator and LLM decision modules, we conduct ablation studies involving two model variations: (1) No Grammar: In this scenario, the grammar translator module is excluded from the framework. LLM agents are prompted to predict candidate performances and select the best candidates directly based on the matrix expression of meta-structures (Han et al., 2020). (2) No Grammar, No LLM Decision: This variation involves the removal of both the grammar translator module and the LLM decision modules from the framework. Candidates are randomly selected to advance to the next generation.
As shown in Figure 8, both ablation models exhibit inferior performance compared to the full model. Specifically, among the three models, No Grammar, No LLM Decision is more prone to getting stuck in local sub-optimal solutions. The LLM selection module facilitates a faster escape from the sub-optimal, showcasing the advantage of knowledge-guided exploration. Ultimately, the full model navigates the search space with significant strides, suggesting that our grammar encoding of the meta-structures contributes to performance enhancement by enabling LLM agents to better grasp the semantic meanings of complex meta-structures. In summary, these results affirm that both designed components contribute considerably to a better meta-structure discovery process.
Figure 8. Comparison of best-performing meta-structure found by each generation.
5. Related Works
5.1. Identifying Meta-structure in HIN
In the past decade, heterogeneous information networks (HIN) gradually gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016). Meta-path, a pre-defined path template of relation sequences, was proposed to measure the similarities between nodes on HIN (Sun et al., 2011). It allows search algorithms like PathSim (Sun et al., 2011) to find peer nodes that are connected by paths with different semantic meanings. The concept of meta-path was later extended to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph. Previous works have found meta-structures useful for boosting machine learning performance on HINs (Jiang et al., 2017). However, early works were based on carefully hand-crafted meta-structures, which heavily relied on domain knowledge. To address this challenge, several more recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021). Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a). We design novel LLM modules for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in several representative datasets.
5.2. Deep Learning on HIN
The success of graph neural networks introduces revolutionary deep learning techniques into HIN modeling (Wu et al., 2020). Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes in HIN via meta-path-guided random walks. Attention mechanism was later introduced to learn more expressive representations for HIN, which were proven effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019). Extensive follow-up research was devoted to designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020). Besides, considerable research efforts were drawn to replace handcrafted meta-structures with automatic search. GEMS (Han et al., 2020) proposed to combine heterogeneous GNN with evolutionary algorithms to identify meta-structures and learn deep neural networks simultaneously. Several deep reinforcement learning models and neural architecture search models are also proposed to jointly optimize the meta-structures with heterogeneous GNN (Ding et al., 2021; Peng et al., 2021; Ning et al., 2022; Li et al., 2023; Wan et al., 2020; Li et al., 2021b). However, previous automatic meta-structure design method solely focused on prediction performance, often yielding complex and difficult to explain meta-structures. To the best of our knowledge, our study is the first to harness the semantic reasoning capability of LLMs for automatic meta-structure design. Our model can discover meta-structures that not only show high prediction performance, but also can be adequately explained with natural languages.
5.3. LLM for Graph Learning
The recent advancement of LLMs shows general capabilities that go beyond natural language tasks. Researchers on graph learning are particularly interested in their ability to graph reasoning tasks (Guo et al., 2023a). For example, there are several recent attempts to use LLMs to enhance node features or employ them as standalone graph predictors (Chen et al., 2023). The research communities are also actively discussing the perspective for developing large graph models (Zhang et al., 2023; Liu et al., 2023). Recent research found LLMs exhibited preliminary ability for graph learning tasks, such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023). Besides, LLMs can also effectively perform reasoning tasks on knowledge graph (Sun et al., 2023). To better align graph problems with LLM, several recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023). On top of these textual encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023). In this paper, we fundamentally extend LLM reasoning to HIN meta-structure discovery. Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLM’s reasoning capability on HINs.
5.4. LLM for Pattern Discoveries
The scaled-up language models have emerged reasoning capability for general tasks (Wei et al., 2022b), including commonsense reasoning, logical reasoning, and mathematics reasoning. Optimized prompting routines like chain-of-thought (CoT) (Wei et al., 2022b) and tree-of-thoughts (ToT) (Yao et al., 2023) can effectively improve the scaling curve of LLM’s reasoning capability. As a result, recent research showed that LLM can be leveraged to identify novel patterns and feasible solutions in a large space. For example, FunSearch was proposed for mathematical algorithms discovery (Romera-Paredes et al., 2023). Besides, previous works also designed novel LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024). In this paper, we propose a novel framework to harness the reasoning power of LLM for meta-structure discovery. Our framework equips LLM with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
6. Conclusion
In this work, we propose a novel framework, ReStruct, that fuses the power of LLMs with evolutionary algorithms to facilitate automatic meta-structure discovery across diverse HINs. Extensive experiments demonstrate that our model excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to state-of-the-art baselines. Notably, a pilot study involving human participants confirms that our model provides a substantial level of interpretability regarding the identified meta-structures. For future work, we will explore the feasibility of finetuning local models to mitigate network communication costs associated with API calls.
Appendix A Statistics of Datasets
Figure 9. Statistics of datasets for node classification
Figure 10. Statistics of datasets for recommendation
Appendix B Questionnaire for Meta-Structure Explainability
Please find the questionnaire in the next few pages.

### v2

@@
 Abstract.
-Heterogeneous information networks (HIN) have gained increasing popularity in the recent decade for their ability to capture the complex relations between nodes of diverse types.
-Meta-structure was proposed to identify the important patterns of relations on HIN, which has been proven effective for extracting rich semantic information from HIN and facilitating graph neural networks to learn expressive representations.
-However, hand-crafted meta-structures pose challenges for scaling up to large and complex HINs, which draw wide research attention for developing automatic meta-structure search algorithms.
-Previous efforts predominantly concentrate on searching for meta-structures that have good empirical prediction performance, overlooking the explainability of discovered meta-structures.
-As a result, they often produce complex meta-structures that are prone to overfitting and incomprehensible to human researchers.
+Heterogeneous information networks (HIN) have gained increasing popularity in recent years for capturing complex relations between diverse types of nodes.
+Meta-structures are proposed as a useful tool to identify the important patterns in HINs, but hand-crafted meta-structures pose significant challenges for scaling up, drawing wide research attention towards developing automatic search algorithms.
+Previous efforts primarily focused on searching for meta-structures with good empirical performance, overlooking the importance of human comprehensibility and generalizability.
 To address this challenge, we draw inspiration from the emergent reasoning abilities of large language models (LLMs).
-In this paper, we propose a novel reasoning meta-structure search (ReStruct) framework that integrates LLM reasoning into the evolutionary procedure for meta-structure search.
-ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate the semantically feasible meta-structures.
+We propose ReStruct, a meta-structure search framework that integrates LLM reasoning into the evolutionary procedure.
+ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate their semantic feasibility.
 Besides, ReStruct also employs performance-oriented evolutionary operations.
-These two competing forces allow ReStruct to jointly optimize for semantic explainability and empirical performance of meta-structures.
-Finally, we also design a differential LLM explainer that can produce natural language explanations for the discovered meta-structures, and refine the explanation by reasoning through the search history.
-Experiments on five representative HIN datasets demonstrate that ReStruct achieve state-of-the-art performance in both node classification and link recommendation tasks.
-Additionally, a survey study involving $73$ graduate students shows that the meta-structures and natural language explanations generated by ReStruct are substantially more comprehensible.
+These two competing forces allow ReStruct to jointly optimize the semantic explainability and empirical performance of meta-structures.
+Furthermore, ReStruct contains a differential LLM explainer to generate and refine natural language explanations for the discovered meta-structures by reasoning through the search history.
+Experiments on eight representative HIN datasets demonstrate that ReStruct achieves state-of-the-art performance in both recommendation and node classification tasks.
+Moreover, a survey study involving $73$ graduate students shows that the discovered meta-structures and generated explanations by ReStruct are substantially more comprehensible.
+Our code and questionnaire are available at https://github.com/LinChen-65/ReStruct.
 Heterogeneous Information Networks, Large Language Models, Graph Neural Networks.
-copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
+copyright: acmcopyrightjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671965isbn: 979-8-4007-0490-1/24/08conference: the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–28, 2024; Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Social networksccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoning
 1. Introduction
-Heterogeneous information networks (HINs) have been proven effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016).
-As a result, HINs are increasingly adopted in a wide range of applications, including social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020).
-To fully exploit the rich semantic information encoded in HINs, researchers proposed to use hand-crafted templates of relation sequences to model the complex proximities on HINs (Sun et al., 2011), which is known as meta-path.
-It was later extended to meta-structure to capture more general interaction patterns than linear paths.
-The recent advancement of heterogeneous GNNs showed that meta-structures can be leveraged to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019), which are crucial for various downstream applications.
-However, the previous works on HINs modeling predominated leverage hand-crafted meta-structures, which heavily rely on the knowledge of domain experts.
-Therefore, they are difficult to scale up to larger and more complex HINs that are vital for real-world applications.
-Motivated by the importance of domain adaptive meta-structure design, numerous recent research efforts are dedicated to designing algorithms for automatic meta-structure design.
-Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search model (Ding et al., 2021) to automatically identify meta-structures that could lead to superior performance in heterogeneous GNNs.
-However, these previous attempts solely focus on the prediction performance of meta-structures, which often lead to highly complex structures that are difficult to explain and susceptible to overfitting.
-Such “meta-structures” deviate from the original inspiration of meta-structures research that aims to extract semantically clear features from HINs (Huang et al., 2016).
-The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to address the challenges of meta-structure discovery.
-The scaled-up versions of LLMs show emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a).
-For example, researchers find chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b).
-Such general-purpose reasoning capability holds huge potential to make sense of the rich semantic information in HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms.
-That is, instead of searching for best-performing meta-structures, the reasoning power of LLM could facilitate the joint optimization for semantic explainability by emulating human reasoners.
-Such reasoning process could effectively prevent the overfitting of identified meta-structures, and produce human understandable knowledge from given HINs.
-In this paper, we propose a novel ReStruct (short for Reasoning meta-Structure search) framework that integrates LLM reasoning into an evolutionary procedure for meta-structure search.
-Specifically, ReStruct employs a grammar translator to encode meta-structures with nested clauses in natural language (see Figure 2).
-It ensures the semantic meaning of meta-structures can be readily comprehended by LLMs as natural language sentences.
-Besides, we define a set of basic operations that can explore the feasible adjacent meta-structures of a given meta-structure, allowing ReStruct to search the adjacent possibilities of promising candidates.
-Different from pure performance-oriented search, ReStruct jointly evaluates the semantic feasibility and empirical performance to determine promising candidates.
-On one hand, ReStruct employs a few-shot LLM predictor to evaluate the semantic similarities between given meta-structure candidates and the searched meta-structures in past experience, and use a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities.
-On the other hand, after the semantically feasible candidates are empirically evaluated with heterogeneous GNNs, ReStruct uses a evolutionary updater adopting the classic elimination-reproduction procedure to mimic natural selection processes that refine meta-structure candidates based on their performances.
+Heterogeneous information networks (HINs) are effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016), leading to their widespread adoption across various applications, such as social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020; Ding et al., 2024).
+To fully exploit the rich semantic information encoded in HINs, researchers have proposed to use meta-paths, which are templates of relation sequences to model the complex proximity on HINs (Sun et al., 2011).
+They were later extended to meta-structures to capture more general interaction patterns beyond linear paths (Huang et al., 2016).
+These meta-structures have been successfully utilized in heterogeneous graph neural networks (GNNs) to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019) for completing downstream tasks.
+However, the reliance on hand-crafted meta-structures, which depend on domain experts’ knowledge, makes it challenging to scale up to larger and more complex HINs that are commonly encountered for real-world applications.
+Driven by the importance of domain adaptation, recent research efforts have been dedicated to developing algorithms for automatic meta-structure search.
+Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search models (Ding et al., 2021) to automatically identify meta-structures that can enhance the performance of heterogeneous GNNs.
+However, these previous attempts primarily focus on the prediction performance of meta-structures, often resulting in highly complex structures that are challenging to interpret and prone to overfitting.
+Such “meta-structures” deviate from the original inspiration of meta-structure research that aims to extract semantically clear features from HINs (Huang et al., 2016).
+The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to tackle the challenges of meta-structure discovery.
+The scaled-up versions of LLMs have exhibited emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a).
+For example, researchers have found that chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b).
+Such a general-purpose reasoning capability holds huge potential for comprehending the rich semantic information and produce human understandable knowledge from given HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms.
+In this paper, we propose a novel framework named ReStruct (short for Reasoning meta-Structure search) that integrates LLM reasoning into an evolutionary procedure for meta-structure search.
+In this framework, we design a grammar translator to encode meta-structures into natural language sentences with nested clauses (see Figure 2), ensuring that their semantic meanings can be readily comprehended by LLMs.
+Besides, we define a set of basic operations to modify a given meta-structure, allowing ReStruct to explore its adjacent possibilities in a valid space.
+Unlike pure performance-oriented search, we anticipate ReStruct to evaluate both semantic feasibility and empirical performance to identify promising candidates.
+To this end, we first design a few-shot LLM predictor to estimate the performance of meta-structure candidates with access to previously evaluated meta-structures from a history pool, followed by a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities.
+After empirically evaluating the chosen candidates with heterogeneous GNNs, we design an evolutionary updater adopting the classic elimination-reproduction procedure to refine meta-structure candidates based on their performances.
 Finally, we design a differential LLM explainer that generates natural language explanations for the discovered meta-structure.
-It employs a chain-of-thought prompting technique to perform structural comprehension and performance attribution step-by-step.
-Such a reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent, not chosen ones.
-We evaluate ReStruct on five representative HIN datasets.
-Experiments show that ReStruct achieve state-of-the-art performance on both node classification and link recommendation tasks.
-Besides, ReStruct can generate natural language explanations as it searches through the solution space.
-To effectively evaluate the explainability of discovered meta-structures, we perform a user survey on $73$ graduate students with domain knowledge of HIN research.
-Survey results show that the meta-structure discovered by our ReStruct method is chosen as the most comprehensible one by 46.6% participants against 3 strong baselines, outperforming the second best baseline by 61.8%.
-More importantly, the natural language explanations generated by our method are preferred by most human participants (77.6% on average) in a head-to-head comparison with baseline methods.
+It employs a chain-of-thought prompting technique to perform step-by-step structural comprehension and performance attribution.
+This reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent yet unchosen ones.
+We evaluate ReStruct on eight representative HIN datasets.
+Experiments show that ReStruct achieves state-of-the-art performance on both recommendation and node classification tasks, and generates meaningful explanations as it searches through the solution space.
+To effectively assess the explainability of the discovered meta-structures, we conduct a user survey on $73$ graduate students with domain knowledge in HIN research.
+According to the survey results, 46.6% of the participants consider the meta-structure discovered by ReStruct as the most comprehensible compared with three strong baselines, outperforming the second best baseline by 61.8%.
+Moreover, the natural language explanations generated by ReStruct are significantly preferred by the majority (77.6% on average) in a head-to-head comparison with baseline methods.
 We summarize our main contributions below:
   • We propose a novel ReStruct framework that integrates LLM reasoning into an evolutionary meta-structure search procedure.
-ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, which coordinates the competing forces of empirical performance-oriented evolutionary updater and semantic similarity-oriented LLM selector.
-ReStruct represents a path-breaking development in meta-structure search algorithms, which could generate meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
+ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, by coordinating the competing forces of an evolutionary updater and a semantic similarity-oriented LLM selector.
+This represents a significant advancement in meta-structure search algorithms, enabling the generation of meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
   • We design a grammar translator to encode meta-structures as natural language sentences, which unleashes the reasoning power of LLMs to make sense of the rich semantic information on HINs.
-On top of this, we design a differential LLM explainer that can generate human comprehensible natural language explanations for discovered meta-structures.
-  • Extensive experiments show that ReStruct achieves state-of-the-art performance on five representative datasets.
-More importantly, a user survey validates that ReStruct substantially outperforms baseline methods in terms of the explainability of discovered meta-structures.
+On top of this, we design a differential LLM explainer that can generate human-comprehensible natural language explanations for discovered meta-structures.
+  • We conduct extensive experiments to reveal ReStruct’s state-of-the-art performance on eight representative datasets.
+Furthermore, we carry out a user survey to validate that ReStruct substantially outperforms baseline methods in terms of the comprehensibility of discovered meta-structures and usefulness of generated explanations.
 2. Preliminaries
@@
 Meta-structure (Huang et al., 2016).
-Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures between the source node type and the target node type.
+Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures beyond linear connections between the source node type and the target node type.
 3. Methods
-In this section, we introduce our proposed methods in detail.
-In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLM’s understanding of its semantic meanings.
-In Section 3.2, we design three basic operations for generating candidate meta-structures.
-In Section 3.3, we design two LLM modules to evaluate and select candidate meta-structures with semantic similarity orientation.
+In this section, we provide a detailed introduction to our proposed methods.
+In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLMs’ understanding of its semantic meanings.
+In Section 3.2, we introduce our design of three basic operations for generating candidate meta-structures.
+In Section 3.3, we design two LLM agents to evaluate and select candidate meta-structures with semantic similarity orientation.
 In Section 3.4, we combine LLM-guided optimization with evolutionary processes to form an effective derivative-free optimization framework.
-Figure 1 shows the overview of our framework.
+The overview of our framework is shown in Figure 1.
 Figure 1. Overview of our proposed ReStruct framework.
+[…]
+Overview of our proposed ReStruct framework.
 3.1. Natural Language Encoding of Meta-Structures
-To facilitate the LLM’s comprehension, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2.
-For any meta-structure, we first traverse the structure to find all possible simple paths from the source node to the target node.
-Each resulting path is equivalent to a meta-path connecting the source and the target, decomposed from the original meta-structure.
-For each path, we encode it into a natural language sentence using nested clauses signified by THAT, which is a commonly-used grammar in English and thus hypothesized to be well-comprehended by the LLM.
-After obtaining the natural language encodings of the decomposed meta-path, i.e., sub-logics, we further combine them using a conjunction word AND to indicate the logical summation effect.
+Previous works represent meta-structures either as matrices or sets of numbers (Han et al., 2020; Ding et al., 2021), which can be challenging to interpret in terms of their semantic meanings.
+As a result, this poses obstacles for LLMs to effectively comprehend such representations.
+To address this limitation and enhance LLMs’ comprehension of meta-structures, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2.
+For a given meta-structure, we begin by traversing its structure to find all possible simple paths connecting the source node to the target node.
+Each resulting path is equivalent to a meta-path decomposed from the original meta-structure.
+Next, we encode each path into a natural language sentence using nested clauses signified by a conjunction word THAT, which is a commonly-used grammar in English and thus expected to be well-comprehended by the LLM.
+In each clause, the central verb connecting two entities is the semantic meaning corresponding to the edge connecting two nodes.
+After obtaining the natural language encodings of the decomposed meta-paths, i.e., sub-logics, we further combine them using another conjunction word AND to convey the logical summation effect.
 Figure 2. Natural language encoding of meta-structures.
+[…]
+Natural language encoding of meta-structures.
 3.2. Basic Operations for Candidate Meta-Structure Generation
-To generate comprehensive candidates for LLM selection and meanwhile ensure validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing Lego.
+To generate comprehensive candidates for LLM selection while ensuring validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing with Lego blocks.
 Examples are shown in Figure 3.
-  • INSERTION. This operation replaces one edge of the original meta-structure with a component.
-  • GRAFTING. This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively.
-  • DELETION. This operation deletes a certain amount of nodes from the original meta-structure, and tries connecting all neighbors back.
+  • INSERTION. This operation replaces one edge of the original meta-structure with a component. It introduces new connections and expands the structure.
+  • GRAFTING.
+This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively.
+It creates branching structures to enhance expressiveness.
+  • DELETION. This operation removes a certain amount of nodes from the original meta-structure, and reconnects the remaining nodes to ensure the structure remains valid.
 In our experiments, we take all meta-paths with no more than $2$ nodes as components for INSERTION, and all meta-paths with no more than $3$ nodes as components for GRAFTING.
 DELETION does not require components as input, as it regards all existing nodes on the original meta-structure as operation candidates.
-Note that using components with more nodes can bring larger exploration space at each step, but may also result in chaos and confusion for the LLM, and thus we leave it as future work to investigate the optimal component settings.
+While using components with more nodes expands the exploration space, it may also introduce complexity and confusion for the LLM.
+We leave it as future work to investigate the optimal component settings.
 Figure 3. Basic operations of exploring adjacent meta-structures.
-3.3. Semantic Similarity-oriented LLM modules for Candidate Selection
-To regulate the explainability of discovered meta-structures, we design two LLM modules that evaluate and select candidate meta-structures in a semantic similarity-oriented manner.
+[…]
+Basic operations of exploring adjacent meta-structures.
+3.3. Semantic Similarity-Oriented LLM Agents for Candidate Selection
+To regulate the explainability of discovered meta-structures, we design two LLM agents that evaluate and select candidate meta-structures in a semantic similarity-oriented manner.
 We illustrate the interaction processes between the main program and the LLM agents in Figure 4.
 3.3.1. Few-Shot LLM Predictor
-Following the execution of the aforementioned basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates.
-This collection comprises meta-structures that have been previously encountered and evaluated in earlier generations, as well as entirely novel ones.
-To inform the decision-making process with insights gleaned from previous evaluations, we design a few-shot LLM predictor to predict the performance $\hat{p}$ of each candidate based on a performance pool that stores the structures and corresponding performances of seen meta-structures.
-Additionally, the few-shot LLM predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate.
-Intuitively, if the few-shot LLM predictor deems a candidate highly similar to a counterpart in the performance pool, it is inclined to assign higher confidence to this prediction.
-This is grounded in the notion that structural similarity is likely indicative of functional similarity.
+After applying the basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates.
+These neighbors can include meta-structures that have been encountered and evaluated in earlier generations, as well as entirely new ones.
+To leverage insights from previous evaluations and guide the decision-making process, we design an LLM agent as a few-shot LLM predictor (abbreviated as “predictor” below).
+This predictor estimates the performance $\hat{p}$ of each candidate through instruction tuning on a small set of structure-performance pairs sampled from a performance pool that records meta-structures in all previous rounds.
+Additionally, the predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate.
+Intuitively, if the predictor considers a candidate to be highly similar to a counterpart in the performance pool, it is likely to predict a similar performance and assign higher confidence to this prediction.
+This is grounded in the understanding that structural similarity often implies functional similarity.
 An illustrative prompt-response round is exemplified in Step 1 of Figure 4.
 3.3.2. Similarity-Oriented LLM Selector
-Receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design an Similarity-oriented LLM selector to make the final decision of picking one single candidate to enter the next generation.
-During this process, the Similarity-oriented LLM selector is expected to jointly consider, and perhaps trade-off between, a meta-structure’s semantic meanings, the expected outcomes of choosing it, as well as the credibility of this outcome expectation.
+Upon receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design another LLM agent, i.e., a similarity-oriented LLM selector (abbreviated as “selector” below), to make the final decision of selecting one single candidate to proceed to the next generation.
+During this process, the selector is expected to consider multiple factors simultaneously, and potentially trade-off between them in order to make the optimal decision (see Appendix B,C).
+These factors include: (1) Semantic meanings, which reflect the relevance and alignment of the meta-structure with the desired objectives and requirements.
+(2) Structural complexities, which indicates the potential risks of overfitting.
+(3) Expected outcomes provided by the few-shot LLM predictor, which indicates the potential benefits from selecting a particular meta-structure in terms of performance improvement.
+(4) Credibility of outcome expectation also provided by the few-shot LLM predictor, which reflects the reliability and trustworthiness of the predictions.
 An illustrative prompt-response round is exemplified in Step 2 of Figure 4.
@@
 Specifically, we maintain a population of $N$ individuals, each representing a distinct meta-structure.
-In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task such as recommendation and node classification.
+In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task.
 After evaluation, the underperforming meta-structures are eliminated from the population.
@@
 Both elimination and reproduction processes mirror natural selection mechanisms that enable species evolution in the wild.
-After getting the modified population, we feed it into the LLM to improve each individual meta-structure.
-Under the evolutionary framework, this can be regarded as a way of targeted mutation, as in this process new nodes and/or edges can be generated and some of the existing nodes and/or edges will be removed.
-The modified population will be further evaluated at the onset of the next generation, forming a loop of derivative-free optimization.
-3.5. Differential LLM Explainer
+After getting the modified population, we feed it into the aforementioned LLM agents to for a new round of individual meta-structure improvement.
+This step can be seen as a way of targeted mutation within the evolutionary framework, as new nodes and/or edges can be generated and some of the existing nodes and/or edges may be removed.
+The modified population will be re-evaluated at the onset of the next generation, forming a loop of derivative-free optimization.
+In summary, by utilizing this evolutionary optimization framework, we can iteratively search for and improve meta-structures without relying on gradient-based optimization methods.
+3.5. Differential LLM Explainer Agent
 One prominent advantage of the LLM lies in its unparalleled ability for natural language generation.
-To harness this ability, we design a differential LLM explainer module to automatically generate human-comprehensible textual explanations elucidating the reasons behind the superior performance of discovered meta-structures.
-We design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) to guide the differential LLM explainer in discerning the critical structural properties that contribute to performance enhancement.
-Specifically, the prompting process for analyzing a given meta-structure $T$ unfolds in the following two steps:
+To harness this ability, we design a differential LLM explainer agent (abbreviated as “explainer” below) to automatically generate human-comprehensible textual explanations that elucidate the reasons behind the superior performance of discovered meta-structures.
+To guide the explainer in discerning the critical structural properties that contribute to performance enhancement, we design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) .
+Specifically, for analyzing a given meta-structure $T$, it unfolds in the following two steps:
 Step 1: Structural Comprehension.
-We first sample a set of $n$ one-step neighbors for $T$, and translate each of them into a natural language sentence according to Method 3.1.
-Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both the original meta-structure and all its sampled neighbors.
-This involves breaking down each of them into meaningful sub-structures and identifying the sub-structure functions.
+We begin by sampling a set of $n$ one-step neighbors for the meta-structure $T$, and translate each of them into a natural language sentence according to Method 3.1.
+Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both $T$ and all its sampled neighbors.
+This process involves breaking down each of them into meaningful sub-structures and identifying the functions os these sub-structures.
 Step 2: Performance Attribution.
 We first perform a quick evaluation of $T$ and each of the sampled neighbors separately by training a GNN with one structure at a time for downstream tasks.
-Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in $T$.
-This involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
-Combination of these two steps empowers the differential LLM explainer to unravel the intricate connections between structural properties and performance, providing a comprehensive understanding of the discovered meta-structures.
-The effectiveness of this module is further justified by a pilot study involving human evaluators, which is elaborated in Section 4.4.2.
+Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in the meta-structure $T$.
+This attribution process involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
+The combination of these two steps empowers the explainer to unravel the intricate connections between structural properties and empirical performance, providing a comprehensive understanding of the discovered meta-structures.
+The effectiveness of this module is further justified by a user study involving human evaluators, which is elaborated in Section 4.4.2.
 Figure 4. Example of LLM prompts and feedbacks.
+[…]
+Example of LLM prompts and feedbacks.
 4. Experiments
@@
 4.1.1. Datasets
-We evaluate our proposed framework ReStruct on two important tasks in heterogeneous information network learning: recommendation and node classification.
-For the recommendation task, the goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., businesses or movies).
-Under this task, we conduct experiments on three widely used real-world datasets: Amazon, Yelp, and Douban  [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding.
-Each dataset includes user ratings of items, which are converted to 0-1 binary labels with a threshold of $2$ to indicate the presence or absence of preference.
-Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into train-validation-test sets with a ratio of 3:1:1.
+We evaluate ReStruct on two important tasks in HIN learning: recommendation and node classification, each with four datasets covering different fields.
+Detailed statistics of all eight datasets can be found in Appendix A.
+In the recommendation task, our goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., items or businesses).
+We conduct experiments on four widely-used real-world datasets: Amazon, Yelp, Douban Movie (abbreviated as ”Douban”), and LastFM [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding.
+For datasets including user ratings of items, the ratings are converted to 0-1 binary labels according to a threshold of $2$.
+A label of 1 indicates the presence of preference, while 0 indicates the absence.
+Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into training-validation-testing sets with a ratio of 3:1:1.
 The other half is reserved for network construction so as to prevent label leakage.
 We take all user-item pairs with the label ’0’ as negative pairs, and also randomly split them into train-validation-test sets to pair each positive pair.
-If the total number of negative pairs is not enough, we randomly sample unconnected items until reaching the desired number.
-We adopt AUC (the area under the ROC curve) as the evaluation metric.
-For the node classification task, the goal is to predict the correct labels of nodes of a given type, such as the genre of a movie.
-Under this task, we conduct experiments on two other widely adopted real-world datasets: ACM and IMDB.
-In these datasets, the classification targeted node type is papers (subjects) for ACM and movies (genres) for IMDB, respectively.
-We follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019).
-We adopt the Macro-F1 score as the evaluation metric.
-Details of all five datasets can be found in Appendix A.
+If the number of negative pairs is insufficient, we randomly sample unconnected items until reaching the desired number.
+The evaluation metric used in these experiments is AUC (Area Under the ROC Curve), which measures the model’s ability to rank positive instances higher than negative instances.
+In the node classification task, our goal is to predict the labels of nodes belonging to a specific type, such as determining the genre of a movie.
+To evaluate the effectiveness of our approach, we perform experiments on four widely-adopted real-world datasets: ACM, IMDB, DBLP, and OAG-NN.
+In these datasets, the classification targets correspond to the subjects of papers in ACM, the genres of movies in IMDB, the research areas of authors in DBLP, and the published venues of papers in OAG-NN, respectively.
+For ACM, IMDB, and DBLP, we follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019).
+For OAG-NN (Hu et al., 2020), we filter the published venues with more than 100 recorded papers, and randomly split the dataset into training-validation-testing sets by 3:1:1.
+The evaluation metric used in these experiments is the Macro-F1 score, which measures the performance of the classification model in terms of precision and recall.
 4.1.2. Baselines
 We compare ReStruct with a set of state-of-the-art baselines. These baselines can be classified into three categories:
-  • Hand-crafted meta-paths: HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism.
-  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework.
-  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which searches meta-structures for recommendation on HINs with a genetic algorithm; (2) DiffMG (Ding et al., 2021), which formalized the problem into a neural architecture search flavor and search for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes with multi-graph search.
+  • Hand-crafted meta-paths: (1) metapath2vec (Dong et al., 2017), which trains a skip-gram model with meta-path guided random walks; (2) HIN2Vec (Fu et al., 2017), which learns latent vectors by jointly training for multiple prediction tasks; (3) HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism; (4) HERec (Shi et al., 2018), which combines random walks with an extended matrix factorization model.
+  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework to search for meta-paths.
+  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which employs a genetic algorithm; (2) DiffMG (Ding et al., 2021), which adopts a neural architecture search manner and searches for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes DiffMG with multi-graph search.
 4.1.3. Hyperparameter Settings
@@
 When modifying each meta-structure, we randomly sample a set of $20$ candidates if there are too many of them from the one-step neighbors.
-When predicting meta-structure performances with the LLM predictor, we randomly sample a set of $30$ evaluation records from the performance pool to enable a few-shot learning paradigm through prompting with examples.
-For the LLM, we use the GPT-4 model by calling the OpenAI API.
-For all baseline models, we follow the implementation released by the authors.
+When predicting meta-structure performances with the few-shot LLM predictor, we randomly sample $30$ records from the performance pool to fuel the few-shot learning paradigm.
+To implement the LLM agents, we use the GPT-4 model by calling the OpenAI API [^2] https://platform.openai.com/docs/models/, while robustness analysis with other LLM models are also carried out (see Section 4.5).
+We employ the DGL implementation of HAN [^3] https://github.com/dmlc/dgl/tree/master/examples/pytorch/han.
+For all the other baseline models, we follow the implementation released by the authors.
 We fix the hidden dimensions to $64$ for all evaluated models, and tune other hyperparameters including learning rate, weight decay, and dropout by referring to the performances on the validation set.
@@
 4.2. Comparison on Recommendation
-In Table 1, we report the experimental results of our proposed framework on the recommendation task compared to baselines.
-First, we observe that models using meta-paths (HAN, RMSHRec) exhibit substantially poorer performances than those using meta-structures (all the rest), confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning.
-Second, ReStruct consistently achieves the best performance across three datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs.
-Specifically, the performance gain over GEMS confirms that LLM-guided “targeted mutation” converges to better solutions that pure random mutation in a classic genetic framework.
-HAN RMSHRec GEMS DiffMG PMMM ReStruct
-Amazon 63.53$\pm$0.24 61.80$\pm$0.15 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
-Yelp 74.54$\pm$0.09 59.39$\pm$0.81 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
-Douban 85.12$\pm$0.12 79.72$\pm$0.32 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
-Table 1. AUC (%) of recommendation on three datasets.
+In Table 1, we report the experimental results of ReStruct on the recommendation task compared to baselines.
+First, we observe that models using meta-paths generally exhibit substantially lower performances than those using meta-structures, confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning.
+Second, ReStruct consistently achieves the best performance across four datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs.
+In particular, the performance gain over GEMS confirms that the LLM-guided “targeted mutation” converges to better solutions than pure random mutation in a classic genetic framework.
+Table 1. AUC (%) of recommendation on four datasets.
+metapath2vec HIN2Vec HAN RMSHRec HERec GEMS DiffMG PMMM ReStruct
+Amazon 56.88$\pm$0.27 58.66$\pm$0.12 60.03$\pm$0.48 61.80$\pm$0.15 70.55$\pm$0.05 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
+Yelp 52.29$\pm$0.59 68.06$\pm$2.31 63.88$\pm$2.32 59.39$\pm$0.81 68.37$\pm$0.34 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
+Douban 52.84$\pm$0.00 85.95$\pm$0.06 63.27$\pm$0.38 79.72$\pm$0.32 92.22$\pm$0.00 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
+LastFM 68.57$\pm$0.10 69.62$\pm$2.00 72.93$\pm$2.42 81.94$\pm$0.22 79.64$\pm$0.06 78.23$\pm$0.08 82.53$\pm$0.11 82.88$\pm$0.07 85.21$\pm$0.09
 Figure 5. Discovered meta-structures on Yelp and the corresponding generated natural language explanations.
+[…]
+Discovered meta-structures on Yelp and the corresponding generated natural language explanations.
 4.3. Comparison on Node Classification
-In Table 2, we report the experimental results of our proposed framework on the node classification task compared to baselines.
-First, HAN and RMSHRec do not achieve desirable performances.
-Specifically, HAN is limited by its heavy reliance on hand-crafted meta-structures, while RMSHRec suffers from high instability presumably due to the limitation accompanied by the reinforcement learning framework.
+In Table 2, we report the experimental results of ReStruct on the node classification task compared to baselines.
+First, metapath2vec, HIN2Vec, and HAN do not achieve desirable performances, mainly due to their heavy reliance on hand-crafted meta-structures.
 Second, DiffMG and PMMM demonstrate improved performances, showcasing the advantage brought by meta-structures over meta-paths, as well as the NAS searching framework.
-Finally, the meta-structure discovered by ReStruct achieves the best performance on both ACM and IMDB datasets.
-HAN RMSHRec DiffMG PMMM ReStruct
-ACM 91.20$\pm$0.25 90.32$\pm$0.74 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
-IMDB 55.09$\pm$0.67 56.72$\pm$2.60 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
-Table 2. Macro F1 scores (%) of node classification on two datasets.
-4.4. Explainability Analysis of Discovered Meta-Structures
+Finally, ReStruct achieves the best performance on ACM, IMDB, and OAG datasets.
+It closely aligns with state-of-the-art results on the DBLP dataset, where Macro F1 scores already exceed $94\%$ – a level challenging to significantly surpass.
+Table 2. Macro F1 scores (%) of node classification on four datasets.
+metapath2vec HIN2Vec HAN DiffMG PMMM ReStruct
+ACM 67.13$\pm$0.50 80.75$\pm$0.77 91.20$\pm$0.25 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
+IMDB 40.82$\pm$1.48 48.16$\pm$0.44 55.09$\pm$0.67 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
+DBLP 89.93$\pm$0.45 90.58$\pm$0.62 92.13$\pm$0.26 94.45$\pm$0.15 94.69$\pm$0.10 94.09$\pm$0.36
+OAG-NN 27.61$\pm$1.65 47.13$\pm$1.14 45.80$\pm$5.84 37.67$\pm$3.13 30.18$\pm$3.87 47.52$\pm$1.56
+4.4. Explainability Analysis
 4.4.1. Visualization of Discovered Structures and LLM-generated Explanation
-In Figure 5, we present $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for the recommendation task, along with a summary text explaining the structural attributes underlying their outstanding performances, which are essentially output from our differential LLM explainer further condensed for clarity.
-We highlight the LLM-identified good (relevant) sub-structures in red, while the bad (distracting) sub-structures in blue.
+Figure 5 showcases $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for recommendation, each with a summary text explaining the structural attributes underlying their outstanding performances.
+These summaries are generated by our differential LLM explainer and further condensed for clarity.
+We highlight the LLM-identified good (relevant) sub-structures in red, and the bad (distracting) sub-structures in blue.
 This visualization facilitates a comprehensive understanding of each structure’s composition.
@@
 These structures, though challenging to handcraft, carry semantically meaningful explanations that remain accessible with textual assistance.
-4.4.2. Pilot Study of Human Evaluation
-We conduct a pilot study with $73$ human participants, more specifically, graduate students with domain knowledge of HIN research, to evaluate the level of explainability provided by our framework compared to baselines.
+4.4.2. User Study for Human Evaluation
+We conduct a user study to evaluate the explainability provided by our framework compared to baselines from a human perspective.
+As our framework targets HIN researchers and engineers as potential users, we recruit $73$ graduate students with domain knowledge of HIN research as our participants.
+To further ensure participants’ solid understanding of the survey’s processes and questions, we provide clear explanations of key concepts such as HIN, meta-path and meta-structure, supported by illustrative examples at the beginning of the survey.
 We structure the study around two sets of questions.
-The first set aims to compare the explainability of the best-performing meta-structure discovered by our model with those found by several baseline models, including GEMS, DiffMG, and PMMM.
-Participants are asked to select one meta-structure from a set of meta-structures, each being the best one found by a specific model, as documented in the original papers (Han et al., 2020; Ding et al., 2021; Li et al., 2023).
-The second set of questions investigates the comprehension gain brought by the textual explanation generated by our differential LLM explainer, as described in Method 3.5, coined the Differential Explanation.
-As a baseline, we include the textual explanation generated by directly prompting an LLM to explain the reasons behind a given meta-structure’s strong performance, without undergoing the two-step prompting process, coined the Non-Differential Explanation.
-For three meta-structures among the best ones discovered by our model on the Yelp HIN, participants are presented with both types of explanations.
-They are then asked to determine which explanation is more helpful in enhancing their (1) understanding of how this meta-structure is constructed, and (2) insights on how to design a better one.
-The complete questionnaire utilized in our study is available in Appendix B.
+The first set of questions is designed to assess the inherent comprehensibility of generated meta-structures without textual explanations.
+To this end, we present the visualizations of the best meta-structures discovered by our model alongside those from three baseline models (GEMS, DiffMG, and PMMM) (Han et al., 2020; Ding et al., 2021; Li et al., 2023), and ask the participants to select the most comprehensible one from their point of view.
+The second set of questions is designed to assess the comprehension gain brought by the textual explanation generated by our differential LLM explainer (see Method 3.5), coined the Differential Explanation.
+As a baseline, we include a Non-Differential Explanation, generated by directly prompting an LLM to explain the reasons behind a meta-structure’s strong performance without undergoing the two-step prompting process.
+For three meta-structures discovered by our model on the Yelp HIN, participants are presented with both types of explanations.
+They are then asked to determine which one is more helpful in enhancing their understanding of how the meta-structure is constructed and gaining insights on how to design a better one.
+By engaging participants in head-to-head comparisons, we aim to gather valuable feedback on the relative helpfulness of each explanation type.
+Before starting the survey, we also carried out a pilot study (Rattray and Jones, 2007) with $5$ participants to ensure the clarity of the questions and visualizations, and none of them expressed confusion or difficulty in understanding these materials.
+The complete questionnaire utilized in our study is available in our GitHub repository.
 Figure 6 shows the result of our first question set.
@@
 Figure 6. Human evaluated explainability of the best meta-structures discovered by different models.
+[…]
+Human evaluated explainability of the best meta-structures discovered by different models.
 Figure 7. Explainability gain brought by LLM-generated differential meta-structure explanation compared to non-differential explanation.
-4.5. Ablation Study
-To assess the efficacy of the designed grammar translator and LLM decision modules, we conduct ablation studies involving two model variations: (1) No Grammar: In this scenario, the grammar translator module is excluded from the framework.
-LLM agents are prompted to predict candidate performances and select the best candidates directly based on the matrix expression of meta-structures (Han et al., 2020).
-(2) No Grammar, No LLM Decision: This variation involves the removal of both the grammar translator module and the LLM decision modules from the framework.
-Candidates are randomly selected to advance to the next generation.
-As shown in Figure 8, both ablation models exhibit inferior performance compared to the full model.
-Specifically, among the three models, No Grammar, No LLM Decision is more prone to getting stuck in local sub-optimal solutions.
-The LLM selection module facilitates a faster escape from the sub-optimal, showcasing the advantage of knowledge-guided exploration.
-Ultimately, the full model navigates the search space with significant strides, suggesting that our grammar encoding of the meta-structures contributes to performance enhancement by enabling LLM agents to better grasp the semantic meanings of complex meta-structures.
-In summary, these results affirm that both designed components contribute considerably to a better meta-structure discovery process.
-Figure 8. Comparison of best-performing meta-structure found by each generation.
+[…]
+Explainability gain brought by LLM-generated differential meta-structure explanation compared to non-differential explanation.
+4.5. Robustness Analysis
+To test the robustness of our method, we first analyze its performance variation under prompt perturbation.
+Following previous studies (Sun et al., 2023a; Moraffah et al., 2024; Zhang et al., 2024), we ask ChatGPT to paraphrase our prompts while maintaining the key module designs in the framework such as grammar translation and LLM-guided performance prediction, and use them to replace the original prompts in the ReStruct framework.
+As shown in Table 3, we find that prompt paraphrasing does not have a clear impact on model performance, as long as the key designs are kept, the importance of which has been validated in the previous section.
+Table 3. Model robustness to prompt perturbation with ChatGPT (taking recommendation on Yelp as an example).
+Index Prompt for Paraphrasing AUC (%)
+0 (Original) - 84.04$\pm$0.24
+[…]
+[…]
+Second, we analyze ReStruct’s robustness by changing the underlying LLM.
+Specifically, we conduct experiments with nine different LLMs from three popular series (GPT, Mistral, and GLM), with different training data, training methods, and parameter sizes.
+As shown in Table 4, the model performances across these LLMs consistently and significantly outperform all baselines, demonstrating the robustness of our framework against LLM versions.
+Table 4. Model performance with different LLM versions (taking recommendation on Yelp as an example).
+Model AUC (%)
+gpt-4-1106-preview (original) 84.04$\pm$0.24
+gpt-3.5-turbo-1106 84.03$\pm$0.11
+mistral-tiny-2312 83.65$\pm$0.17
+mistral-small-2312 83.89$\pm$0.09
+mistral-small-2402 83.96$\pm$0.12
+mistral-medium-2312 84.04$\pm$0.16
+mistral-large-2402 83.87$\pm$0.13
+glm-3-turbo 84.12$\pm$0.06
+glm-4 83.76$\pm$0.07
 5. Related Works
-5.1. Identifying Meta-structure in HIN
-In the past decade, heterogeneous information networks (HIN) gradually gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016).
+5.1. Identifying Meta-structures on HINs
+Over the past decade, HINs have gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016).
 Meta-path, a pre-defined path template of relation sequences, was proposed to measure the similarities between nodes on HIN (Sun et al., 2011).
 It allows search algorithms like PathSim (Sun et al., 2011) to find peer nodes that are connected by paths with different semantic meanings.
-The concept of meta-path was later extended to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph.
+The concept of meta-path was later extended beyond the linear relationship to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph.
 Previous works have found meta-structures useful for boosting machine learning performance on HINs (Jiang et al., 2017).
-However, early works were based on carefully hand-crafted meta-structures, which heavily relied on domain knowledge.
-To address this challenge, several more recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021).
-Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a).
-We design novel LLM modules for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in several representative datasets.
-5.2. Deep Learning on HIN
+However, early works were based on carefully hand-crafted meta-structures, which heavily relied on experts’ domain knowledge.
+To address this challenge, several recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021).
+Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a) for this task.
+We design novel LLM agents for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in eight representative datasets.
+5.2. Deep Learning on HINs
 The success of graph neural networks introduces revolutionary deep learning techniques into HIN modeling (Wu et al., 2020).
-Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes in HIN via meta-path-guided random walks.
-Attention mechanism was later introduced to learn more expressive representations for HIN, which were proven effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019).
-Extensive follow-up research was devoted to designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020).
+Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes via meta-path-guided random walks.
+Attention mechanism was later introduced to learn more expressive representations for HINs, proving effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019).
+Subsequent research endeavors focused on designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020).
 Besides, considerable research efforts were drawn to replace handcrafted meta-structures with automatic search.
@@
 5.3. LLM for Graph Learning
-The recent advancement of LLMs shows general capabilities that go beyond natural language tasks.
-Researchers on graph learning are particularly interested in their ability to graph reasoning tasks (Guo et al., 2023a).
-For example, there are several recent attempts to use LLMs to enhance node features or employ them as standalone graph predictors (Chen et al., 2023).
-The research communities are also actively discussing the perspective for developing large graph models (Zhang et al., 2023; Liu et al., 2023).
-Recent research found LLMs exhibited preliminary ability for graph learning tasks, such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023).
-Besides, LLMs can also effectively perform reasoning tasks on knowledge graph (Sun et al., 2023).
-To better align graph problems with LLM, several recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023).
-On top of these textual encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023).
+LLMs have demonstrated general capabilities beyond natural language tasks, attracting graph learning researchers who are particularly interested in leveraging their ability for graph reasoning tasks (Guo et al., 2023a).
+Several attempts have been made to enhance node features using LLMs or employ them as standalone graph predictors (Chen et al., 2023).
+The research community is also actively discussing the perspective of developing large graph models (Zhang et al., 2023; Liu et al., 2023).
+Recent research has found that LLMs exhibit certain ability for graph tasks such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023).
+Besides, LLMs can effectively perform reasoning on knowledge graphs (Sun et al., 2023b).
+To better align graph problems with LLMs, recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023).
+With these encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023).
 In this paper, we fundamentally extend LLM reasoning to HIN meta-structure discovery.
-Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLM’s reasoning capability on HINs.
+Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLMs’ reasoning capability on HINs.
 5.4. LLM for Pattern Discoveries
 The scaled-up language models have emerged reasoning capability for general tasks (Wei et al., 2022b), including commonsense reasoning, logical reasoning, and mathematics reasoning.
-Optimized prompting routines like chain-of-thought (CoT) (Wei et al., 2022b) and tree-of-thoughts (ToT) (Yao et al., 2023) can effectively improve the scaling curve of LLM’s reasoning capability.
-As a result, recent research showed that LLM can be leveraged to identify novel patterns and feasible solutions in a large space.
-For example, FunSearch was proposed for mathematical algorithms discovery (Romera-Paredes et al., 2023).
-Besides, previous works also designed novel LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024).
+With the help of optimized prompting routines such as chain-of-thought (CoT) (Wei et al., 2022b) prompting and tree-of-thoughts (ToT) (Yao et al., 2023) prompting, the scaling curve of LLMs’ reasoning capability can be further effectively improved.
+As a result, recent research has shown that LLMs can be leveraged to identify novel patterns and feasible solutions in large problem spaces.
+For example, FunSearch was proposed to discover algorithm programs for solving mathematical problems (Romera-Paredes et al., 2023).
+Previous works also designed LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024).
 In this paper, we propose a novel framework to harness the reasoning power of LLM for meta-structure discovery.
-Our framework equips LLM with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
+Our framework equips LLMs with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
 6. Conclusion
 In this work, we propose a novel framework, ReStruct, that fuses the power of LLMs with evolutionary algorithms to facilitate automatic meta-structure discovery across diverse HINs.
-Extensive experiments demonstrate that our model excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to state-of-the-art baselines.
-Notably, a pilot study involving human participants confirms that our model provides a substantial level of interpretability regarding the identified meta-structures.
+On both recommendation and node classification tasks, extensive experiments demonstrate that ReStruct excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to a set of state-of-the-art baselines.
+Notably, a user study involving human participants confirms that ReStruct substantially outperforms baseline methods in terms of the comprehensibility of discovered meta-structures and usefulness of generated explanations.
 For future work, we will explore the feasibility of finetuning local models to mitigate network communication costs associated with API calls.
+7. Acknowledgement
+This work was supported in part by the National Key Research and Development Program of China under Grant 23IAA02114, the Guangzhou Municipal Nansha District Science and Technology Bureau under Contract No.2022ZD012, and the National Natural Science Foundation of China under Grant U23B2030 and Grant U22B2057.
 Appendix A Statistics of Datasets
-Figure 9. Statistics of datasets for node classification
-Figure 10. Statistics of datasets for recommendation
-Appendix B Questionnaire for Meta-Structure Explainability
-Please find the questionnaire in the next few pages.
+Table 5. Statistics of datasets for node classification.
+Dataset ACM IMDB DBLP OAG-NN
+Node
+types
+Author (A)
+Paper (P)
+Subject (S)
+Movie (M)
+Actor (A)
+Director (D)
+Author (A)
+Paper (P)
+Conference (C)
+Paper (P)
+Author (A)
+Affiliation (I)
+Field (F)
+Edge
+types
+A-P, P-A,
+P-S, S-P
+M-D, D-M,
+M-A, A-M
+A-P, P-A,
+P-C, C-P
+P-P,
+P-A, A-P,
+P-F, F-P,
+A-I, I-A
+# Nodes 8,994 12,624 18,405 64,203
+# Edges 25,922 37,288 67,946 403,974
+# Classes 3 3 4 8
+# Training 600 300 800 2,334
+# Validation 300 300 400 778
+# Testing 2,125 2,339 2,857 778
+Table 6. Statistics of datasets for recommendation.
+Dataset Relations (S-T) # S # T # S-T
+Yelp User-Business (U-B) 16,239 14,284 84,993
+User-User (U-U) 16,239 16,239 158,590
+User-Compliment (U-O) 16,239 11 76,875
+Business-Category (B-A) 14,284 511 40,009
+Business-City (B-I) 14,284 47 14,267
+Douban User-Movie (U-M) 13,367 12,677 500,515
+User-User (U-U) 13,367 13,367 4,085
+User-Group (U-G) 13,367 2,753 570,047
+Movie-Actor (M-A) 12,677 6,311 33,587
+Movie-Director (M-D) 12,677 2,449 11,276
+Movie-Type (M-T) 1,2677 38 27,668
+Amazon User-Item (U-I) 6,170 2,753 86,191
+Item-View (I-V) 2,753 3,857 5,694
+Item-Category (I-C) 2,753 22 5,508
+Item-Brand (I-B) 2,753 334 2,753
+LastFM User-Artist (U-A) 1,892 17,632 46,417
+User-User (U-U) 1,892 1,892 25,434
+Artist-Tag (A-T) 17,632 9,718 108,437
+Appendix B Emergence of Occam’s Razor Phenomenon
+In our user study, we demonstrate that ReStruct finds meta-structures that are more comprehensible to human researchers (Figure 6), implying a tendency to avoid overcomplicated meta-structures.
+To provide further evidence, we add an experiment on the Yelp recommendation task, asking the LLM to select between meta-structure pairs with near-equal performance on the validation set but varying structural complexity (#nodes and #edges), which are visualized in Figure 8.
+As shown in Table 7, LLM consistently prefers meta-structures with lower structural complexity, citing reasons such as ”simplicity,” ”generalizability,” and even invoking ”Occam’s razor”.
+Therefore, LLMs possess human-like preferences for simpler and more comprehensible meta-structures that mitigates overfitting.
+Figure 8. Visualization of meta-structure pairs for comparison.
+[…]
+Visualization of meta-structure pairs for comparison.
+Table 7. LLM selection between meta-structure pairs with near-equal validation performances: three examples.
+Pair Chosen Index Reasons
+1 (Fig.
+8A) 1 “- Complexity vs.
+Performance: The second meta-structure has a more complex path, suggesting potentially richer semantics being captured.
+However, if both meta-structures show almost equal performance, the additional complexity might not translate into a significant performance gain.
+In such cases, it is generally better to opt for the simpler model (Occam’s razor), which is why Meta-structure 1 is preferred.
+- Overfitting Risks: The more complex second meta-structure might be more prone to overfitting, as it captures a more specific set of interactions.
+The simpler Meta-structure 1 might generalize better to unseen data.
+- Interpretability: Meta-structure 1 is easier to interpret because it has a shorter path and fewer node interactions.
+This can be beneficial for understanding why certain recommendations are being made, which is important for transparency and trust in the recommendation system.
+- Computational Efficiency: With shorter paths and fewer relations, Meta-structure 1 is likely to be more computationally efficient for both training and inference, which can be important in practical, large-scale applications.
+- Meta-path Relevance: The paths in Meta-structure 1 adequately capture the relevant user-business interactions via a business’s location.
+Adding more nodes and relations in the second meta-structure might not contribute much to capturing the essential features for the recommendation task.
+In conclusion, the first meta-structure is selected as it achieves a good balance between simplicity, performance, and interpretable results while potentially being more computationally efficient and generalizable.”
+2 (Fig.
+8B) 1 “- Path Simplicity: Meta-structure 1 offers a more straightforward path from ’U’ (User) to ’B’ (Local Business) through ’A’ (Business Category) and ’I’ (Business City), with minimal backtracking or redundant connections.
+The path follows a logical progression from user preference to business through the context of category and location.
+- Computational Efficiency: The simpler meta-structure 1 should, in theory, be computationally less intensive because it involves fewer steps and connections.
+This could be advantageous when scaling the recommendation system and processing a large dataset.
+- Interpretability: Meta-structure 1 is easier to interpret.
+In the context of a recommendation system, a clear logical flow from user to business via category and city is likely to align with a user’s thought process when seeking recommendations, enhancing the user experience.
+- Relevance of Categories and Location: By going through ’A’ (Business Category) and ’I’ (Business City), Meta-structure 1 reflects important aspects of businesses that users consider while making choices—what kind of service or product a business offers (category) and where it is located (city).
+While Meta-structure 2 includes a potentially rich and complex connection pattern that might capture more nuanced relationships, the balance between complexity and utility must be considered, particularly if both structures have shown similar performance on the validation set.
+Given these considerations, the simpler, more interpretable, and likely more efficient Meta-structure 1 would likely be the optimal choice for the recommendation model.”
+3 (Fig.
+8C) 1 “- **Simplicity and Interpretability**: The first meta-structure is less complex as it contains a shorter sequence of nodes, potentially making it easier to interpret and explain how the recommendation is generated.
+A simpler model is often preferred when performances are almost equal because it can be more transparent, and it can generalize better, avoiding overfitting the data.
+- **Computational Efficiency**: Fewer node hops suggest lower computational costs in terms of executing the meta-path.
+This can be important for scaling the recommendation model to a large number of users or in an environment where computational resources are constrained.
+- **Redundancy Avoidance**: The second meta-structure involves a repeated sequence of ’Business Category’ to ’Local Business’ (’A’ to ’B’) hops, which may introduce redundancy that does not necessarily contribute to predictive performance.
+Such over-specified meta-structures can sometimes lead to overfitting, where the performance on the validation set is good, but the generalization to unseen data may suffer.
+- **Logical Coherence**: The first meta-structure provides a more direct flow of information from ’User’ to ’Local Business’ through their categories and mutual connections to other users.
+This logical flow is more coherent compared to the second meta-structure, which seems to loop back through ’Business Category’ without a clear additional benefit.
+- **Real-World Significance**: From a real-world perspective, the first meta-structure captures a user’s preference through the categories of businesses they have visited and the influence of their friends’ preferences for the same business categories.
+This could be a robust basis for recommendations without the need for the extra category-business-category loop in the second meta-structure.
+Given these reasons, meta-structure 1 seems to be the optimal choice for improving the recommendation model on this Yelp HIN.”
+Appendix C Evidence of opportunity-risk tradeoff during candidate selection
+To check whether there is an overreliance on confidence values during candidate selection, we add an experiment to compare the confidence values between chosen and unchosen meta-structure candidates.
+As shown in Figure 9, ReStruct does not adhere strictly to the candidates with the highest confidence values.
+Take Experiment #14 as an example: although Candidate 00 and $5$ have higher confidence values, ReStruct decides to choose Candidate $3$, likely due to its higher predicted performance via effective semantic analysis.
+This highlights ReStruct’s attempts to balance between opportunities and risks, when predicted performances and confidence values become trade-offs.
+Figure 9. Comparison of confidence values between chosen and unchosen candidates.
+[…]
+Comparison of confidence values between chosen and unchosen candidates.